A sample of water requires 238.46 J to be heated from 21.83 oC to 38.48 oC. What is the mass of the water?
1
Expert's answer
2019-10-18T08:06:46-0400

m=Q/c*(t1-t2)=238.46J/4200 (J/kg°C)*16.65°C=0.0034kg=3.4g